OpenStudy (redbirdd):

Solve for the roots in the following equation. Hint: Factor both quadratic expressions. (x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0

OpenStudy (will.h):

Do we factor both expressions separately or do we simplify the expression and then factor? Which one do they want

OpenStudy (mikey):

2x * x = 2x^2

OpenStudy (mikey):

did that help?

OpenStudy (freckles):

factor x^4+5x^2-36 and then factor 2x^ 2+9x-5 I like to find two numbers that multiply to be a*c and add up to b then factor by grouping (of course you can do short cut way if a=1)

OpenStudy (mathmale):

You have (x^4 + 5x^2 - 36)(2x^2 + 9x - 5) = 0, and are to find the roots. Factor the polynomial within each set of parentheses: Factor (x^4 + 5x^2 - 36) as far as possible, and then factor (2x^2 + 9x - 5) as far as possible. Then, set each of your factors = to 0 and find the solution(s). List all of the solutions as your final answer.

OpenStudy (mathmate):

Hint: 1. When there are only even powers in a polynomial to be factored, replace \(x^{2n}\) by \(y^n\) and proceed to factor the reduced polynomial. Example: Factor \(a^4-b^2\) substitute y=a^2, then \(a^4-b^2=y^2-b^2=(y+b)(y-b)=(a^2+b)(a^2-b)\) 2. Use the usual rules to factorize each quadratic. Do not multiply out. For great examples, read through all simple, hard and weird cases at the following link: http://www.purplemath.com/modules/factquad.htm